Define simple harmonic motion and give the expressions for the velocity and acceleration of such motion at the different points along it's path
1
Expert's answer
2021-08-13T10:20:25-0400

Simple harmonic motion is a kind of motion that occurs periodically. Such motion can be described with sine or cosine function.

Velocity at any point:


"v(t)=v_\text{max}\cos(\omega t),"

Acceleration:


"a(t)=-a_\text{max}\sin(\omega t)=\\-\omega v_\text{max}\sin(\omega t)."